### v3

# Aligning Large Language Models to a Domain-specific Graph Database for NL2GQL

Abstract
Graph Databases (Graph DB) find extensive application across diverse domains such as finance, social networks, and medicine. Yet, the translation of Natural Language (NL) into the Graph Query Language (GQL), referred to as NL2GQL, poses significant challenges owing to its intricate and specialized nature. Some approaches have sought to utilize Large Language Models (LLMs) to address analogous tasks like text2SQL. Nonetheless, in the realm of NL2GQL tasks tailored to a particular domain, the absence of domain-specific NL-GQL data pairs adds complexity to aligning LLMs with the graph DB. To tackle this challenge, we present a well-defined pipeline. Initially, we utilize ChatGPT to generate NL-GQL data pairs, leveraging the provided graph DB with self-instruction. Subsequently, we employ the generated data to fine-tune LLMs, ensuring alignment between LLMs and the graph DB. Moreover, we find the importance of relevant schema in efficiently generating accurate GQLs. Thus, we introduce a method to extract relevant schema as the input context. We evaluate our method using two carefully constructed datasets derived from graph DBs in the finance and medicine domains, named FinGQL and MediGQL. Experimental results reveal that our approach significantly outperforms a set of baseline methods, with improvements of 5.90 and 6.36 absolute points on EM, and 6.00 and 7.09 absolute points on EX for FinGQL and MediGQL, respectively.
Aligning Large Language Models to a Domain-specific Graph Database for NL2GQL
Yuanyuan Liang1, Keren Tan1, Tingyu Xie2, Wenbiao Tao1 , Siyuan Wang1 Yunshi Lan1thanks: *Corresponding author,Weining Qian1 1 East China Normal University, 2 Zhejiang University {leonyuany,tankeren1020,wbtao,sywang}@stu.ecnu.edu.cn tingyuxie@zju.edu.cn, {yslan,wnqian}@dase.ecnu.edu.cn
1 Introduction
Presently, leading Graph Databases (Graph DBs) such as Neo4j, NebulaGraph, and JanusGraph are extensively utilized for storing graph data Wang et al. (2020); Lopes et al. (2023). These Graph DBs have broad applications across various domains, including finance, healthcare, social networks, and e-commerce. However, the nuanced and complex syntax of the Graph Query Language (GQL) presents hurdles for general users endeavoring to convert Natural Language (NL) into GQL. Figure 1 illustrates a NL2GQL example. A GQL consists of keywords MATCH, WHERE, and RETURN leading the retrieval functions towards a graph DB.
Several factors combine to render NL2GQL an exceedingly challenging task. Primarily, the inherent complexity of NL understanding poses initial hurdles due to its intricate nature. Additionally, the intricate syntax of GQL further compounds the difficulty Monteiro et al. (2023). Furthermore, the diversity of terminology and conventions within domain-specific graph data adds another layer of complexity to the translation process.
Figure 1: An instance of the NL2GQL task in the financial domain, which entails the conversion of NL into GQL tailored for a specific graph DB.
While the NL2GQL task is both highly significant and extremely challenging, recent research on this topic is remarkably scarce. The predominant research emphasis is on the conversion of NL into Cypher, which is an efficient GQL aligned to the Nebula Graph. Guo et al. (2022) released a general-domain Text-to-Cypher dataset and explored the application of a basic Seq2Seq framework for this task. Nevertheless, the accuracy of their proposed baseline method is notably low, barely surpassing a meager $2\%$. Zhou et al. (2023) developed a hybrid approach combining smaller and larger foundation models to improve NL2GQL accuracy in general domain. Considering the connection between NL2GQL and text2SQL tasks, Zhao et al. (2024) developed a SQL2Cypher algorithm capable of mapping SQL queries to Cypher queries. However, the relationship between schema items relies heavily on the domain knowledge, which cannot be easily generalized from general domain. Above approaches are hard to be applied to NL2GQL over a domain-specific graph DB.
In the last year, Large Language Models (LLMs) have transitioned from emerging technologies to widely used applications Wei et al. (2022a); Zhao et al. (2023); Kaddour et al. (2023); Xi et al. (2023). Concurrently, a body of research substantiates the robust capacity of LLMs in transforming NL into logical forms Luo et al. (2023); Meyer et al. (2023); Li et al. (2023) with a lack of training data. Although these methods have achieved certain results, their methods are tailored to Knowledge Graphs (KGs), which are stored in Resource Description Framework (RDF) Pan (2009) and have distinct schema structure to graph DBs, there is still a significant gap for LLMs application in NL2GQL.
In order to improve the performance of NL2GQL tasks on a specific graph DB without any label data, we propose a well-defined pipeline, as depicted in Figure 2. First, we generate NL-GQL pairs as instructions based on a given Graph DB for fine-tuning LLMs. Specifically, starting from the Graph DB, we collect a set of canonical NL-GQL template pairs. Subsequently, we employ a two-step self-instruct Wang et al. (2022) method to extend these initial pairs, generating a comprehensive collection of NL-GQL template pairs, and further being verified via Chain-of-thought Wei et al. (2022c). This ensures broad query coverage and diverse expressions while maintaining consistency between NL template and GQL template. The resulting NL-GQL template pairs undergo graph DB grounding to get complete NL-GQL data pairs. Secondly, we align a foundation LLM to the graph DB with LoRA Hu et al. (2021). During inference, we organize the schema that is relevant to NL for prompting the aligned LLMs. We evaluate the performance of our aligned on the graph DBs in finance and medicine domains. We select traditional fine-tuned methods and In-context Learning (ICL) methods as baseline methods. The experimental results demonstrate that our method outperforms all of the baseline methods by a noticeable margin.
We summarize our contributions as follows:
  • We introduced a pipeline for aligning LLMs with a domain-specific Graph DB, enabling the seamless incorporation of pertinent information from the graph database and its stored data into LLMs.
  • To enhance the accuracy and inference speed of LLM on NL2GQL tasks, we propose a method that extracts a portion of schema from the graph database schema, which is relevant to NL, to serve as input context.
  • Experimental results showcase the superior effectiveness of our proposed approach compared to the baseline.
Figure 2: Overview of our method.
2 Related Work
2.1 NL2GQL
NL2GQL is an emerging task that requires a system to convert a natural language question into formats compatible with graph DB. The system should have strong capabilities of question understanding and GQL parsing. Initial endeavors focused on direct translations from natural language to queries, primarily relying on hand-crafted rules Zhao et al. (2022). More recent studies shift towards data-driven approaches. By training a sequence-to-sequence model on NL2GQL datasets, a system can learn the transition from NL to GQL well Guo et al. (2022). Our study demonstrates innovation in solving NL2SQL over a domain-specific graph DB and aligning LLMs to such a graph DB.
We also notice that text2SQL and KGQA (question answering over knowledge graphs) are analogous tasks to NL2GQL, where the query language are compatible with traditional DBs and knowledge bases, respectively. Even though prior studies have developed a number of text2SQL as well as KGQA datasets and end-to-end systems Gao et al. (2023a); Chang and Fosler-Lussier (2023); Gu et al. (2023); Saxena et al. (2020); Jiang and Usbeck (2022), they cannot be directly transferred to NL2GQL. NL2GQL proves its differences in complex graph modality, variance query types and distinct graph query languages Zhou et al. (2023). While our approach aligns with NL2SQL in terms of the methodological flow, there are substantial disparities in solving NL2GQL over a domain-specific graph DB.
2.2 Fusing LLMs and Databases
LLMs have shown its powerful capabilities in a diverse range of NLP tasks Min et al. (2023); Kaddour et al. (2023); Ferrara (2023). Incorporating LLMs with databases is an important research question Pan et al. (2024); Luo et al. (2023); Gao et al. (2023b); Pourreza and Rafiei (2023); Zhou et al. (2023), which benefits to improve the efficiency of structured data processing and ease the hallucination issues of LLMs. The integration of LLMs and structured DB is primarily approached through two main paradigms. The first paradigm extracts the information from databases to enhance the contexts for prompting LLMs in generation. For instance, techniques like Retrieval Augmented Generation (RAG) Gao et al. (2023c) can be employed to retrieve and query relevant data from databases, thereby enhancing the generated results Kang et al. (2023a); Wang et al. (2023); Modarressi et al. (2023); Luo et al. (2023); Yang et al. (2023b); Gu et al. (2023); Zhou et al. (2023) during inference.
The second paradigm injects the knowledge of databases into LLMs by fine-tuning De Cao et al. (2021); Lin et al. (2021); Sen et al. (2023); Kang et al. (2023b). A key aspect of this approach is to make controllable edit to the parameters in LLMs. Our approach take advantages of both paradigm. We tuned a foundation LLM to align it to knowledge of a graph DB on the one hand. On the other hand, we elaborately extract relevant schema from the graph DB for prompting, which deeply fuses LLMs and databases together.
3 Preliminary
3.1 Graph Databases
A graph database $\mathcal{G}$ is a relational database, which is stored in the form of $\mathcal{G}=\{(s,r,o)|s,o\in\mathcal{E},r\in\mathcal{V})\}$. Here $\mathcal{E}$, $\mathcal{V}$ denotes the vertex set and the edge set, respectively. Graph database is a property graph, where every vertex and edge can be defined with labels and have any number of properties (aka attributesBesta et al. (2023). Each graph database is associated with a schema $\mathcal{S}$.
3.2 NL2GQL with LLMs
When it comes a natural language query $Q$ inquiring about the information in $\mathcal{G}$, NL2GQL task requires a system to transfer $Q$ into a graph database language query $L$ (e.g., Cypher, nGQL, and Gremlin) which can be executed over $\mathcal{G}$ to retrieve the predicted answer. In this paper, we treat an LLM as an NL2GQL system to map $Q$ to $L$ with the consideration of pre-defined $\mathcal{G}$.
4 Our Method
In order to align LLMs to a Graph DB and achieve NL2GQL with high accuracy, we propose a pipeline as illustrated in Figure 2. Initially, we generate NL-GQL pairs as instruction for fine-tuning the LLM. Specifically, starting from the Graph DB, we collect a set of canonical NL-GQL templates annotated with GQL. Then we extend these seed pairs to a set of NL-GQL pairs with a wide query coverage and diverse expressions via two-step data augmentation based on ChatGPT. These generated NL-GQL pairs are further grounded and verified. Next, we fine-tune LLMs via LoRA with the generated instruction such that the tuned LLM will be aligned with the knowledge in the graph DB. For inference, to bound the LLMs with the graph DB schema, we organize the relevant schema to the query as the context of the prompt and request the LLM to produce GQL using the identified schema.
4.1 Generating NL-GQL Pairs as Instructions
It is non-trivial to construct NL-GQL data pairs deriving from a given $\mathcal{G}$. Different from the majority of existing QA over DB challenges like Spider Yu et al. (2018), GrailQA Gu et al. (2021) and SpCQL Guo et al. (2022), where the narrative questions are associated with a database in general domain. $\mathcal{G}$ is in specific domain such that it is not applicable to employ the above datasets as initialization. In addition, it is impossible to be scaled up if we solely rely on human labor to annotate the NL-GQL pairs in a specific domain. Therefore, it is vital to develop a labor-saving pipeline to generate high-quality NL-GQL pairs that are able to cover a wide range of query types on a specific domain.
Therefore, we employ ChatGPT to play an important role in the pipeline of NL-GQL pairs generation. With the awareness that ChatGPT have lack of the specific knowledge of $\mathcal{G}$, a NL-GQL template collector module is developed to collect the canonical NL-GQL templates which decouple the redundant schema items from the GQL. Then, a two-step instruction generator with self-instruct module is implemented. Its first step strengthens the guidance to ChatGPT on $\mathcal{S}$ via providing the schema of graph DB as well as canonical NL-GQL templates as context for prompting. In its second step, the outputs are verified via GQL2NL with reasoning the intermediate steps. Furthermore, we continue to improve the quality and coverage of instructions by self-supervising the instruction tuning process until the desired number of NL-GQL pairs is obtained. Eventually, we ground the NL-GQL pairs via a graph DB grounding module.
NL-GQL template collector. To ensure the diversity and comprehensiveness of generated NL-GQL on different graph DB, we formulate an NL-GQL template as a narrative question with masked schema items (i.e., entities and properties) followed by a GQL skeleton, where the entity labels and properties of the node are masked out. An example is shown below:
  • What is [entity]’s [property]?
  • MATCH (s:node{name:‘[entity]’})
    RETURN s.node.[property] .
where we employ the placeholder [entity] to denote the named entities and [property] to denote the property of an entity in both the NL and GQL.
We define $8$ types of the canonical NL-GQL templates covering queries about entity property, numerical sorting, attribute comparison and so on [^1] The detailed query types and their corresponding templates can be found in Appendix A. For each NL type, we manually annotate $K$ data pairs and include them into a data pool $\mathcal{D}$. As we can see, the NL-GQL templates are able to decouple the schema-related information for the queries while keeping the main semantics and syntax, which can be adapted to different domains.
[Schema of Graph DB]:
Nodes
   {’tag’: ’stock’, ’properties’: [(’code’, ’string’), …}
Edges
   {’edge’: ’has_stock_data’,’start_tag’:’stock’,
       ’end_tag’:’stock_data’,’properties’: []}
[Task Description]: Please generate a new NL-GQL
template data pair of the same type as the provided NL
based on the given Graph database schema and NL-GQL
demonstrations. In the …
[NL-GQL]:
What is [entity]’s stock code?
MATCH (s:stock{name:’[entity]’})
RETURN s.stock.code
Table 1: The prompt for NL-GQL data pair generation.
Two-step instruction generator with self-instruct. The first step aims to generate candidate NL-GQL pairs with $\mathcal{S}$ as well as $\mathcal{D}$ and using ChatGPT. We design a unified prompt but adjust to the different types of query by changing instructions for each category. As shown in Table 1, each prompt consists of the following components:
  • [Schema of Graph DB]: According to Gao et al. (2023b), including the DB schema in the prompt helps the generation of SQL that fits the DB. Therefore, our prompt used for generating NL-GQL data pairs consists of the entire schema of the graph DB. Specifically, we organize the schema by flattening the nodes set into a sequence, wherein each node type is represented as a key accompanied by its corresponding properties. Similarly, each edge is defined with its start and end node types, along with any associated properties, as depicted in Table 1.
  • [Task Description]: It is a textual sequence describing the query type of pairs to be generated, we pre-define the description in Appendix C.
  • [NL-GQL]: We sample $K$ examples of a certain type from $\mathcal{D}$ and randomly sample properties from the schema to instantiate the [property] variable, resulting in different demonstrations.
In this case, we can prompt LLMs to generate NL-GQL pairs that are related to the graph DB. For simplicity, we denote this procedure as:
$Q^{\prime},L^{\prime}=\text{NL-GQL-Gen}(\mathcal{S},\mathcal{D}).$
The second step takes $L^{\prime}$ as the input and aims to verify NL with GQL generated from the first step. Several studies Wei et al. (2022c); Kojima et al. (2022); Zhou et al. (2022); Diao et al. (2023) have established the efficacy of utilizing CoT Wei et al. (2022b) to yield more accurate answers by generating intermediate reasoning steps. In this step, we apply CoT to conduct GQL-to-NL, the goal of which lies in verifying the quality of the generated NL given a GQL. The prompt format closely resembles that of Table 1, with the task-adjusted description and replacement of [NL-GQL] with [CoT-based GQL2NL] demonstrations:
  • [CoT-based GQL2NL]: We take the $K$ examples in the first step but present the transition from GQL to NL in the format of a explicit reasoning chain. Specifically, we break down the GQL into individual keywords function clauses based on keywords such as MATCH, WHERE and RETURN. For example, the given GQL is MATCH (s: stock{name: ‘[entity]}) RETURN s.stock.code. Its CoT process is as follows: MATCH (s: stock{name: ‘[entity]} identifies the entity with the stock name [entity]. RETURN s.stock.code returns the code of the stock. Combining these two parts, so the output is: "What is [entity]’s code?"
Therefore, we can obtain an alternative question by prompting LLMs with the given GQL. For simplicity, we denote this procedure as:
$Q^{{}^{\prime\prime}}=\text{GQL2NL-Gen}(\mathcal{S},L^{\prime}).$
To ensure the consistency of $L$ and $Q$ across all data entries, we compare encode them via the sequence matching encoder and compare their similarities. If it falls below a certain threshold, we consider the $Q^{\prime}$ and $L^{\prime}$ to be inconsistent, leading to the discarding of the respective data entry. Otherwise, we include the NL-GQL pairs into data pool $Q^{\prime},L^{\prime}\in\mathcal{D}$.
The two step generation results in high-quality NL-GQL pairs. To enhance the diversity of NL-GQL pairs, we conduct self-instruct Wang et al. (2022) for the generation procedure, where the demonstrations used for the next iteration are extracted from $\mathcal{D}$. This iterative process continues until we generate enough NL-GQL pairs.
Graph DB grounding. In the previous steps, all generated NL and GQL pairs in $\mathcal{D}$ are templates with [entity] placeholders. Then, we replace these placeholders with named entities from the graph DB to obtain instantiation of NL as well as GQL. The GQLs are executed over the associated graph DB to check if they are semantically and syntactically correct to retrieve the answers. Only the executable NL-GQL pairs will saved. As a result, we have a high-quality data pool in specific domain with each entry containing an NL, GQL, template NL, template GQL, and the corresponding executed answers.
4.2 Aligning LLMs to a Graph DB with LoRA
After we generate a set of NL-GQL pairs, we can fine-tune an existing foundation LLM to align with the graph DB. To avoid extensive parameter tuning, we apply Parameter Efficient Fine-Tuning (PEFT) techniques, which is able to minimize fine-tuning parameters and optimize memory usage while maintaining the performance Liu et al. (2022); Hu et al. (2021). We choose LoRA Hu et al. (2021) for fine-tuning the selected LLMs due to its wide application. Specifically, we feed the NL-GQL pairs obtained from Section 4.1 into the LLMs. For each pair, we extract the schema items in the GQL as the context to the NL as an instruction, then supervised fine-tuning calculates the cross-entropy loss over the ground truth GQL. In this way, the fine-tuned LLMs learn the schema of the domain-specific graph DB and understand the transition from the NL to the GQL.
4.3 Organizing Schema for Prompting
Regarding inference for NL2GQL tasks, we prompt the tuned LLM to generate $L$ given $Q$ directly. Similar attempts have been carried on Text2SQL by representing a DB schema with a linear “code” sequence and appending the sequence as the context for prompting Nan et al. (2023), we prepend the schema of the graph DB as context but differ in the way of formulation. Specifically, we extract the key schema items that are relevant to the NL rather than the entire schema as the context. To this end, a label linker is used to extract the named entities from $Q$ and link them to the labels in the graph DB. Then, a LLM prompt organizes the extracted labels as the context and generates the main body of the GQL. Eventually, a join-table linker further infers the joint tables based on the extracted labels and append it to the main body to form a prompt.
Label linker. To accurately identify named entities in NL and link them to the labels in the graph DB, we first collect all the labels from the graph DB and store them in a dictionary, where the labels are saved as the key and their corresponding properties are saved as the value. When there comes a $Q$, we tokenize the sentence and employ an off-the-shelf named entity tagger to identify the named entities within it. Then, we map these named entities to possible keys in the dictionary via string match. This step may result in multiple ambiguous labels. For example, the named entity “Sam” could be linked to either a fund manager or a chairman. We disambiguate the named entities by choosing the labels in a GQL from $\mathcal{D}$, the NL of which has the highest similarity to $Q$.
Join-table linker. Since many questions inquire about information that requires reasoning across multiple tables, but not all table information is explicitly mentioned in the question, it is necessary to perform table linking with the linked labels. We formulate the problem as searching for the shortest path through the tables where the identified labels are located. We utilize the A* algorithm Hart et al. (1968) to address the shortest path problem. This results in the shortest chain linking all the identified labels, which serves as the MATCH keyword function in a GQL.
[Relevant Schema of Graph DB]:
Nodes
    {’tag’: ’stock’, ’properties’: [(’code’, ’string’), …}
Edges
    {’edge’: ’has_stock_data’,’start_tag’:’stock’,
        ’end_tag’:’stock_data’,’properties’: []}
[Task Description]: You are a graph database expert.
Please write the corresponding graph query language based
on the relevant schema and natural language.
[NL]: Which stocks have a minimum price below 20 and
an opening price below 30 today?
Table 2: The prompt for fine-tuning LLMs.
The MATCH keyword function of inferred GQL is verified by the outputs of joint-table linker to form an executable GQL.
LLM prompt. After acquiring the relevant schema, incorporate the task description and input natural language to compose a comprehensive prompt, as shown in Table 2.
5 Experiments
5.1 Experimental Setup
Datasets FinGQL MediGQL
train dev test train dev test
Graph DB info
#Nodes $668508/12$ types $44656/8$ types
#Edges $730583/14$ types $312159/12$ types
#Properties $62$ types $14$ types
Dataset info
#Nodes (Avg.) $1.81$ $1.78$ $1.81$ $1.76$ $1.84$ $1.73$
#Edge (Avg.) $0.81$ $0.78$ $0.81$ $0.86$ $0.90$ $0.83$
#Unique Template $1434$ $229$ $419$ $1907$ $293$ $590$
#NL-GQL $7000$ $1000$ $2000$ $6000$ $850$ $1700$
Table 3: Statistics of the generated datasets.
Evaluation Set Construction. To verify the way we fine-tune a foundation LLM and prompt the tuned LLM for solving NL2GQL over a domain-specific graph DB, we choose StockKG [^2] We collect financial data from open stock exchange websites: http://www.sse.com.cn/market/price/report/, https://www.szse.cn/market/overview/index.html, https://sc.hkex.com.hk/TuniS/www.hkex.com.hk/Market-Data/Securities-Prices/Equities and DiseaseKG [^3] http://openkg.cn/dataset/disease-information as the graph DB in finance and medicine domains, respectively. We follow the pipeline description in Section 4.1 to generate a number of NL-GQL pairs. We randomly partitioned the data pool into training, development, and test sets with ratios of $7:1:2$, respectively. The correctness of test sets are further verified manually. These two data sets, namely FinGQL and MediGQL, have statistics shown in Table 3 and can be leveraged to support future research on NL2GQL over domain-specific graph DB [^4] More details about the presented datasets are provided in Appendix B.. We will release the datasets upon acceptance.
     Method      Backbones      FinGQL      MediGQL
     EM      EX      EM      EX
     Fine-Tuning      mbart-large      69.55      53.45      64.92      54.56
     mt5-large      73.75      67.75      79.69      65.04
     In-Context Learning      Baichuan2-13B-chat      18.20      14.90      23.54      18.24
     Chatglm3-6B      16.10      17.75      18.78      16.13
     Qwen-14B-Chat      21.10      22.75      24.24      21.07
     ChatGPT-3.5      40.65      40.10      49.38      41.38
     Ours      Baichuan2-13B-Chat      78.10      71.25      84.54      71.38
     Chatglm3-6B      77.30      68.95      83.81      70.51
     Qwen-14B-Chat      79.65      73.75      86.05      72.13
Table 4: Comparison between our method and the baseline method. The bold numbers indicate the best results.
Evaluation Metrics. We follow the evaluation methods of text2SQL, which include exact-set-match accuracy (EM) and execution accuracy (EXYu et al. (2018). EM measures the consistency of the individual components split by keywords within the predicted query and its corresponding ground truth, whereas EX compares whether the execution results on the database are consistent between the two.
Baseline Methods. To validate and compare the effectiveness of our method, we selected two categories of methods as baselines:
  • Fine-tuning methods. Instead of constructing instructions for fine-tuning, we employ the traditional approach Lu et al. (2022); Guo et al. (2022) to fine-tune a pre-trained model, where the NL from NL-GQL data pairs serves as input, and the GQL serves as the label for training.
  • In-Context Learning methods. We follow the RAG paradigm to include the schema of graph DB as the context and sample $K$ examples with large semantic similarities to the NL as the demonstrations for prompting LLMs during inference Dong et al. (2023); Gao et al. (2024).
Implementation Details. In selecting the backbone LLMs, we prioritize LLMs compatible with both Chinese and English due to the language of employed Graph DB. As a result, we choose Baichuan2-13B-Chat Yang et al. (2023a), Chatglm3-6B Zeng et al. (2022), and Qwen-14B-Chat Bai et al. (2023) as the foundation LLMs to be aligned. Regarding fine-tuning methods, we select mbart-large Tang et al. (2020) and mt5-large Xue et al. (2021) as the pre-trained models due to their wide application in multilingual scenarios. Regarding the sequence matching for demonstration selection, we utilized the all-MiniLM-L6-v2 checkpoint from the Sentence-Transformers library to encode a sequence. The threshold for verifying the NLs from two steps in NL-GQL pair generation is set as $0.8$. All the number of demonstrations $K$ are set as $8$. Our experiments were implemented on an A800 GPU.
5.2 Main Results
Table 4 illustrates a performance comparison between our method and the baseline methods. From the experimental results, we can draw the following conclusions:
a) Our approach demonstrates a significant improvement over the best baseline methods. Specifically, on the FinCQL dataset, it outperforms them by 5.90% on the EM metric and by 6.00% on the EX metric. Similarly, on the MediCQL dataset, it surpasses them by 6.36% on the EM metric and by 7.09% on the EX metric.
b) Backbone model is crucial for the NL2GQL task. The mt5-large model consistently outperforms the mbart-large model. Additionally, both the ICL method and our approach demonstrate that the Qwen-14B-chat model yields significantly better results compared to other open-source LLMs. Furthermore, within the ICL method, ChatGPT-3.5 exhibits notably superior performance compared to several other LLMs, although it only achieves roughly half the accuracy of our approach.
c) It’s evident that a simplistic ICL method isn’t suitable for NL2GQL. This may be attributed to the lack of domain knowledge from the data within the graph DB. To improve the effectiveness of the ICL method, further research is needed in prompt design to better guide LLMs in extracting domain knowledge.
Furthermore, it is evident from the result table that EM values generally exceed those of EX. EM serves as a more forgiving evaluation metric, while EX provides a more precise assessment by directly comparing GQL execution outcomes.
5.3 Comparison with Domain-Specific LLMs
One of the reasons why the ICL performs poorly on the NL2GQL task is its insufficient understanding of the domain-specific knowledge. Therefore, we are interested in exploring whether Domain-Specific LLMs can perform better. We experiment on two well-known medical-domain LLMs, HuatuoGPT2-7B Chen et al. (2023a) and BianQue-2 Chen et al. (2023b), and the results are in Table 5. As shown in the table, compared to the general-domain Chatglm3-6B and Qwen-14B-Chat, the domain-specific HuatuoGPT2-7B and BianQue-2 LLMs perform very poorly. It is presumably because they focused too much on domain-specific knowledge and tasks, neglecting general knowledge and tasks. During the domain-specific supervised fine-tuning, the ability of language conversion of LLMs is largely compromised, resulting in poor performance in the NL2GQL task.
Domain Model EM EX
Medical domain HuatuoGPT2-7B 0.059 0.059
BianQue-2 0.18 0.18
General domain Chatglm3-6B 16.10 17.75
Qwen-14B-Chat 21.10 22.75
Table 5: The comparison of performance between medical domain LLMs and general domain LLMs using the ICL method on the MediGQL dataset.
5.4 Further Analysis
Efficiency Analysis. To observe the inference speed of our method, we conducted experiments on the test set of FinGQL and recorded its execution time. The results are shown in Table 6. Based on Table 6 and Table 8, it is evident that our strategy of incorporating the relevant Chinese schema leads to a substantial enhancement in effectiveness, accompanied by a slight increase in inference time. However, adding the full schema will significantly increase the inference time, and including unrelated distracting information may also lead to a decrease in accuracy.
Method Execute time (Sec.)
Without Schema 0.77
Full Schema 1.32
Full Chinese Schema 1.52
Relevant Schema 0.87
Relevant Chinese Schema (Ours) 0.88
Table 6: Comparison of average inference times for various schema-related methods on the FinGQL dataset, measured in seconds.
Breakdown Analysis. For a more detailed analysis of our method’s performance across various query types, we segmented the execution results by query type and tabulated the statistics, as presented in Table 7. From the table, it can be observed that the accuracy is relatively low for queries of Yes/No question type, Edge property type, and Attribute comparison type. These three types of queries all involve complex computational problems, so there may be some challenges in generating GQL with accurate calculation logic.
Query type EM EX
Entity property 81.87 80.63
Numerical sorting 84.88 84.88
Relationship inference 84.18 84.48
Yes/No question 69.40 68.87
Relationship filtering 79.55 77.27
Attribute comparison 76.01 78.50
Edge property 70.15 71.61
String filtering 88.89 85.19
Total 79.65 73.75
Table 7: Performance of our method on various types of queries in the FinGQL dataset.
5.5 Ablation Study
We conducted an ablation study to evaluate the effectiveness of components within our method, and the results are presented in Table 8. The results indicate that removing the Chinese explanations of fields in the schema leads to a deterioration in performance, suggesting their contribution to addressing the Chinese NL2GQL task. Moreover, without any schema leads to a more significant performance drop, highlighting the beneficial role of schema as context in addressing the NL2GQL task.
Method EM EX
Relevant Chinese Schema (Ours) 86.05 72.13
Relevant Schema 85.12 $\downarrow$(0.9) 71.81 $\downarrow$(0.3)
Without Schema 83.46 $\downarrow$(2.6) 69.92 $\downarrow$(2.2)
Full Schema 85.29 $\downarrow$(0.8) 71.98 $\downarrow$(0.1)
Full Chinese Schema 85.55 $\downarrow$(0.5) 71.34 $\downarrow$(0.8)
Table 8: Ablation study of our method on MediGQL. The green downward arrow denotes a decrease, and the green number in parentheses indicates the precise decrease value.
6 Conclusion
In this paper, we present a pipeline for addressing the NL2GQL task on a given graph DB. Firstly, we construct an NL2GQL dataset based on the graph database. Then, we fine-tune LLMs using the constructed dataset to align LLMs with the graph database. Additionally, to further enhance the NL2GQL capability of LLMs, we propose a method that extracts schema relevant to NL as input context to guide LLM. Experimental results demonstrate that our approach significantly outperforms baseline methods.
Limitations
Although our work is intended to be language-agnostic, it is regrettable that our experiments currently only involve Chinese data, as we only have access to domain graph data available in Chinese.
Appendix A Query Type
Based on the logical structure between nodes, edges, and properties that may be involved in queries, we summarized the queries and classified them into 8 query types, which are presented in Table 10.
Appendix B Dataset Presentation
The FinGQL dataset comprises a total of 10,000 records, each consisting of an NL along with its corresponding GQL, a template NL along with its corresponding template GQL, the corresponding GQL Answer, and the Query Type ID. Furthermore, Nodes and Edges are extracted from the GQL to provide Relevant Schema during training. Similarly, the MedGQL dataset consists of 8,550 records.
An example from the FinGQL dataset is as follows:
  • NL:
    {CJK*}UTF8gbsn 钢铁行业中,有哪些股票的开盘价大于1?
    (In the steel industry, which stocks have opening prices greater than 1?)
  • GQL:
    MATCH (t:trade{name:"{CJK*}UTF8gbsn钢铁"(steel)})
    <-[bt:belong_to]-(s:stock)-[hsd:has_stock_data]->(sd:stock_data) WHERE sd.stock_data.opening_price > 1 RETURN s.stock.name
  • template NL:
    {CJK*}UTF8gbsn [t]行业中,有哪些股票的开盘价大于[m]?
    (In the [t] industry, which stocks have opening prices greater than [m]?)
  • template GQL:
    MATCH (t:trade{name:’[t]’})
    <-[bt:belong_to]-(s:stock)-[hsd:has_stock_data]->(sd:stock_data) WHERE sd.stock_data.opening_price > [m] RETURN s.stock.name
  • Query Type ID: 6
  • Nodes: ["stock", "stock_data"]
  • Edges: ["has_stock_data", "belong_to"]
Appendix C NL-GQL Data Pairs Generation Task Description
We aim to guide ChatGPT in generating NL-GQL template data pairs of the desired query type. Merely including schema and demonstrations in the prompt is not sufficient. Detailed descriptions of the desired query types and explanations of the specific constraints associated with each type are also necessary. Our task description follows the format below:
Please generate a new NL-GQL template data pair of the same type as the provided NL based on the given Graph database schema and NL-GQL demonstrations. In the Graph database schema section, Nodes contain information about all entities, including node tags and their corresponding properties. Edges contain information about all relationships, including the head and tail nodes of the edge and any edge properties. In the NL-GQL demonstrations, named entities are replaced with placeholders. For example, ’stock’ is represented by ’[s]’, ’chairman’ by ’[c]’, ’trade’ by ’[t]’…(Description of all replacement characters for nodes and edges in the schema is provided below. Due to space constraints, only a portion is written.) The query type is { query type }, in this type { query type description}.
In the above task description, replace "{query type}" and its corresponding "{query type description}" with the type of NL to be generated. The query type descriptions we summarized are shown in Table 11.
Appendix D Ablation Study on FinGQL
We present Table 9 to showcase ablation studies on FinGQL datasets. Additionally, we recognize the importance of relevant Chinese schema in assisting LLMs to generate more accurate GQL. It is worth noting that we did not conduct experiments with the full schema approach due to the 2048 input length limit of Qwen-14B-Chat. Adding the full schema would exceed this limit. This underscores the importance of the strategy to extract the relevant schema of NL in our method.
Method EM EX
Relevant Chinese Schema 79.65 73.75
Relevant Schema 77.85 $\downarrow$(1.8) 71.85 $\downarrow$(1.9)
Without Schema 77.35 $\downarrow$(2.3) 71.55 $\downarrow$(2.2)
Table 9: Ablation study of our method on FinGQL. The green downward arrow denotes a decrease, and the green number in parentheses indicates the precise decrease value.
Query type Type#1 Entity property Type#2 Numerical sorting
NL template What is [entity]’s [property]?
The [property2] of node with the highest
[property1]?
GQL template
MATCH (s:node{name:’[entity]’})
RETURN s.node.[property]
MATCH (s:node) WITH s.node.[property2] AS n1,
s.node.[property1] AS n2 ORDER BY n2 DESC
LIMIT 1 RETURN n1
Query type Type#3 Relationship inference Type#4 Yes/No question
NL template The [property] of node1 rel2 by [entity]?
Is the [property] of the [entity]
greater than [mount]?
GQL template
MATCH (s1:node1)<-[:rel1]-(s2:node2)
<-[:rel2]-(s3:node3{name: ’[entity]’})
RETURN s1.node1.[property]
MATCH (s:node{name: ’[entity]’})
RETURN s.node.[property] >[mount]
Query type Type#5 Relationship filtering Type#6 Attribute comparison
NL template
The [property] of node2 that rel [entity1]
and [entity2]?
How much does the [property] of [entity1]
differ from that of [entity2]?
GQL template
MATCH (s1:node1{name: ’[entity1]’})<-[:rel]
-(s2:node2)-[:rel]->(s3:node1{name: ’[entity2]’})
RETURN s2:node2.[property]
MATCH (s1:node{name: ’[entity1]’})
WITH s1.node.[property] AS t1
MATCH (s2:node{name: ’[entity2]’})
WITH ABS(s2.node.[property] - t1)
as abs_diff RETURN abs_diff
Query type Type#7 Edge property Type#8 String filtering
NL template
The [property] of node2 that rel [entity] with
[r_property] less than [m]?
The [property] of the node containing
the [string]?
GQL template
MATCH (s1:node1{name: ’[entity]’})-[r:rel]
->(s2:node2) WHERE r.[r_property] <[m]
RETURN s2.node2.[property]
MATCH (s:node) WHERE s.node.[property]
CONTAINS ’[string]’ RETURN s.node.[property]
Table 10: 8 canonical query types along with their corresponding NL and GQL templates.
Query type Query type descriptions
Type#1 Entity property The generated NL and GQL are both related to simple property queries of nodes and do not involve any computations.
Type#2 Numerical sorting The generated NL and GQL are both related to sorting concepts and numerical values.
Type#3 Relationship inference The generated NL and GQL are both related to reasoning about relationships between nodes and edges.
Type#4 Yes/No question The generated NL and GQL both pertain to ’whether’ or ’existence’ related queries.
Type#5 Relationship filtering The generated NL and GQL are both related to filtering relationships on multi-hop subgraphs associated with nodes.
Type#6 Attribute comparison The generated NL and GQL are both related to aggregate calculations of node properties.
Type#7 Edge property The generated NL and GQL are both related to comparing and aggregating calculations of edge properties.
Type#8 String filtering The generated NL and GQL are both related to string matching.
Table 11: Description of each query type.